OpenStudy (anonymous):

4x2 + 25x + 6 D.) (2x + 6)(2x + 1) C.) (2x + 3)(2x + 2) B.) (4x + 6)(x + 1) A.) (4x + 1)(x + 6)

OpenStudy (souvik):

find 2 numbers which adds to 25 and multiplied to 4*6=24

OpenStudy (anonymous):

I don't understand how to work it at all

OpenStudy (austinl):

You need to find two numbers that when multiplied give you 4. Every option has that. So the next step is you need to find two numbers that when multiplied and added to another give 25.

OpenStudy (anonymous):

but how do I even know what numbers to use???

OpenStudy (austinl):

4*1=4 4*6=24 +1=25 A lot of it is educated guesses, and practice.

OpenStudy (austinl):

So, in conclusion.... (A) is the correct choice.

OpenStudy (anonymous):

thanks!

OpenStudy (jdoe0001):

your numbers will be the factors of "6" in this case that is the factors of the constant factors of "6" are 3, 2, 1, 6 and the factors of 4 that is the leading term coefficient factors of 4 are 2, 2, 1, 4

OpenStudy (austinl):

Also, for problems similar to this, you could also FOIL out the possibilities and get your answer.

OpenStudy (anonymous):

how do you use FOIL I never understood how

OpenStudy (austinl):

FOIL= First Outer Inner Last

OpenStudy (austinl):

Lets use option D as an example. (2x + 6)(2x + 1) You would multiply 2x & 2x to get 4x^2 Then you would multiply the outer terms, 2x & 1 to get 2x Then the inner terms, 6 & 2x to get 12x And Last, 6 & 1 to get 6 \[4x^2+2x+12x+6\] Condense to get \[4x^2+14x+6\]

OpenStudy (anonymous):

o ok that helps thanks!

OpenStudy (jdoe0001):

so, why does A work? well, let's see $$ (4x + 1)(x + 6)\\ 4x \times x = \color{red}{4x^2}\\ (4x \times 6) + (1 \times x) \implies 24x + x = \color{red}{25x}\\ +1 \times +6 = \color{red}{+6}\\ \color{red}{4x^2+25x+6} $$

OpenStudy (anonymous):

that makes a lot more sense thanks!